OpenStudy (anonymous):

Explain how to write an equation of a line in standard form when two points on the line are given.

OpenStudy (anonymous):

The general equation describing a line is y = m*x + b if you have got two points, solve first for one of the variables m or b and the for the other

OpenStudy (anonymous):

Example: you have the points (2, 4) and (3, 8), then you plug any in the equation, say (2, 4), and you get: 4 = m*2 + b, so b = 4 - 2m now you do the other: 8 = m*3 + b, and you know b, so: 8 = 3m + (4 - 2m) = 4 - m, therefore m = -4 nnow you put that in the b equation and get: b = 4 - 2m = 4 - 2*-4 = 12 So our line is y = -4x + 12 :)